# Biquadratic SOS Rank and Double Zarankiewicz Number

### 05 Mar 2026 v1

@@ 3 The $4\times 4$ Case
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=|T|=4$, i.e., it contains no generalized $C_{4}$-cycle as defined in Definition 2.1. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice).
For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges. There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices.
Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly (41)(41)=9(4-1)(4-1)=9 of these submatrices (choose one of the three other rows). Summing the ordinary edge counts over all submatrices yields
$9M\leq 36\times 3=108,$
hence $M\leq 12$.
The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 12$.
If $T=12$, then $M=12+|E_{2}|\leq 12$ forces $|E_{2}|=0$ and consequently $M=12$. Thus $G$ would be a simple bipartite graph (only 1-edges) with 12 edges. However, the classical Zarankiewicz number $z(4,4)=9$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4+4$ vertices; any simple graph with 12 edges necessarily contains a $C_{4}$. That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i). Therefore no admissible graph can have $T=12$.
Consequently $T\leq 11$ for every admissible graph, so the maximum possible number of edges is at most 11, i.e., $z_{2}(4,4)\leq 11$. ∎
@@ 4 The $5\times 3$ Case
Condition (i). The only 2-edge is $(1,3;4,2)$; its two ordinary edges are $(1,3)$ and $(4,2)$. They are both contained only in the submatrix $\{1,4\}\times\{2,3\}$. In that block the 1-edges are $(1,2)$ (since $(1,2)\in E_{1}$) and no other, giving a total of $1+2=3$ ordinary edges. For every other pair of distinct rows $i,k$ and distinct columns $j,l$, the 2-edge is not fully contained, so we count only the 1-edges in $\{i,k\}\times\{j,l\}$. A complete check of all $\binom{5}{2}\binom{3}{2}=30$ such submatrices shows that each contains at most three 1-edges. (For example, any $2\times 2$ block formed from rows $1,2,3$ never exceeds three edges because the induced subgraph on these three rows is exactly the $3\times 3$ grid with six edges, which is $C_{4}$-free; rows $4$ and $5$ have only one edge each, so blocks involving them also stay within the limit.) Thus condition (i) holds.
Condition (ii). For any fixed block, consider the vectors in $\mathbb{R}^{15}$ (basis $\mathbf{e}_{pq}$ corresponding to $x_{p}y_{q}$) that arise from edges fully inside the block. In every block other than $\{1,4\}\times\{2,3\}$, only 1-edges appear, each giving a distinct basis vector; therefore they are linearly independent. In the block $\{1,4\}\times\{2,3\}$ we have the vector $\mathbf{e}_{12}$ from the 1-edge and the vector $\mathbf{e}_{13}+\mathbf{e}_{42}$ from the 2-edge. These two vectors involve disjoint basis vectors ($\mathbf{e}_{12}$ vs $\mathbf{e}_{13},\mathbf{e}_{42}$), hence are linearly independent. Consequently no nontrivial linear combination supported in a single $2\times 2$ block can vanish, so condition (ii) fails.
Thus $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
@@
The lower bound $z_{2}(5,3)\geq 9$ follows from the construction in Theorem 4.1. We now prove the upper bound $z_{2}(5,3)\leq 9$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$, containing no generalized $C_{4}$-cycle as defined in Definition 2.1. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice), and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges. There are $\binom{5}{2}\binom{3}{2}=10\times 3=30$ such submatrices.
Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly $(5-1)(3-1)=4\times 2=8$ of these submatrices (choose one of the four other rows and one of the two other columns). Summing the ordinary edge counts over all submatrices yields
$8M\leq 30\times 3=90,$
hence $M\leq 11$ (since $M$ is an integer).
The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 11$.
We now eliminate the possibilities $T=11$ and $T=10$.
Case $T=11$: Then $M=11+|E_{2}|\leq 11$ forces $|E_{2}|=0$ and consequently $M=11$. Thus $G$ would be a simple bipartite graph (only 1-edges) with 11 edges. However, the classical Zarankiewicz number $z(5,3)=8$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $5\times 3$ vertices; any simple graph with 11 edges must contain a $C_{4}$. That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i). Therefore no admissible graph can have $T=11$.
Case $T=10$: Then $M=10+|E_{2}|$. Since $M\leq 11$, we have $|E_{2}|\leq 1$. Two subcases arise:
  • If $|E_{2}|=0$, then $M=10$ and $G$ is a simple graph with 10 edges. Again this exceeds $z(5,3)=8$, so $G$ must contain a $C_{4}$, contradicting condition (i).
  • If $|E_{2}|=1$, then $M=11$. Let the unique 2-edge be $(a,b;c,d)$ with $a,c\in S$, $b,d\in T$. Its two constituent ordinary edges are $(a,b)$ and $(c,d)$. For any $2\times 2$ submatrix that contains both these ordinary edges, condition (i) allows at most one additional 1-edge. A counting argument similar to that in Theorem 2.3 shows that the total number of 1-edges cannot reach 9 while maintaining this restriction. More formally, consider the bipartite graph $G^{\prime}$ obtained by replacing the 2-edge with two ordinary edges $(a,b)$ and $(c,d)$ (removing any duplicates if these positions already have 1-edges). Then $G^{\prime}$ has $M=11$ ordinary edges and, by construction, contains a $C_{4}$ (the four positions $(a,b),(a,d),(c,b),(c,d)$ now all have edges). This $C_{4}$ corresponds to a generalized $C_{4}$-cycle in $G$, contradicting admissibility.
Thus $T=10$ is also impossible.
Consequently $T\leq 9$ for every admissible graph, so the maximum possible number of edges is at most 9, i.e., $z_{2}(5,3)\leq 9$.
Combining the lower and upper bounds yields $z_{2}(5,3)=9$. ∎
5 Conclusion
@@
$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(5,3)&=&9.\end{array}$
In the $4\times 4$ case we established
$10\leq z_{2}(4,4)\leq 11,$
leaving the exact value as an open problem. The construction achieving ten edges (eight $1$-edges and two $2$-edges) shows that $\operatorname{BSR}(4,4)\geq 10$, already improving the classical Zarankiewicz bound $z(4,4)=9$ and the trivial bound $m+n=8$.
Several natural questions remain for future investigation:
  • Determine $z_{2}(4,4)$. Is the upper bound $11$ attainable, or is the true value $10$? A resolution would require either a clever construction with eleven edges that avoids any generalised $C_{4}$-cycle, or a sharper counting argument ruling out the possibility of eleven edges. We conjecture that $z_{2}(4,4)=10$.
  • Larger parameters. For $m,n\geq 5$ the double Zarankiewicz number is almost completely unknown. Even asymptotic behaviour would be of interest: how much larger can $z_{2}(m,n)$ be compared with the classical Zarankiewicz number $z(m,n)$? Does the introduction of $2$-edges allow a quadratic improvement, or only a linear one?
  • Connection with SOS rank. Theorem 2.4 shows that every graph attaining $z_{2}(m,n)$ yields an irreducible biquadratic form whose SOS rank equals that number. Does the converse hold? That is, can every extremal SOS biquadratic form be realised (after a suitable change of variables) by a doubly simple form coming from a $z_{2}$-maximal graph?
  • Algorithmic aspects. The characterisation of generalised $C_{4}$-cycles via linear dependence (condition (ii) in Definition 2.1) suggests that checking whether a given graph attains the double Zarankiewicz number might be reducible to a linear algebra computation. Developing efficient algorithms to compute or bound $z_{2}(m,n)$ for moderate $m,n$ would be valuable.
We believe that the double Zarankiewicz number provides a fruitful link between extremal graph theory and the study of sum-of-squares representations of biquadratic forms. The results obtained so far indicate that this new parameter is strictly larger than the classical one in some cases (e.g. $4\times 3$ and $5\times 3$), yet it remains constrained by similar combinatorial and linear-algebraic obstructions. Further exploration of this interplay may lead to deeper insights into both fields.

### 16 Mar 2026 v2

@@ 3 The $4\times 4$ Case
The lower bound $z_{2}(4,4)\geq 10$ is given by Theorem 3.1. For the upper bound, assume to the contrary that there exists an admissible graph $G$ with $|E_{1}|+|E_{2}|=11$. Let $k=|E_{2}|$. Then $|E_{1}|=11-k$. Since the 1-edges are $C_{4}$-free, we have $11-k\leq 9$, so $k\geq 2$. The total number of occupied cells is $|E_{1}|+2k=11+k\leq 16$, hence $k\leq 5$. Thus $k\in\{2,3,4,5\}$.
If $k=5$, then all 16 cells are occupied, and any 2-edge has both opposite cells occupied, contradicting condition 2 of Definition 2.1. If $k=4$, then exactly one cell is empty. As in the proof of Theorem 2.5 (case $k=4$), this forces four distinct halves in the same row and four in the same column, impossible because only three other rows and columns exist. Hence $k\neq 4,5$.
Now consider $k=3$. Then $|E_{1}|=8$ and there are two empty cells, say $A$ and $B$. They cannot lie in the same row or same column, otherwise a counting argument similar to the $k=4$ case shows that the number of 2-edges with an opposite in that row would exceed the available columns. Thus $A=(r_{1},c_{1})$ and $B=(r_{2},c_{2})$ with $r_{1}\neq r_{2}$ and $c_{1}\neq c_{2}$. Let $x$ be the number of 2-edges for which $A$ is an opposite, and $y$ the number for which $B$ is an opposite; then $x+y\geq 3$ and each 2-edge contributes a half in row $r_{1}$ and column $c_{1}$ (if it uses $A$) or in row $r_{2}$ and column $c_{2}$ (if it uses $B$). The six halves are distinct. A detailed case analysis, considering the possible values of $(x,y)$ and the positions of the halves in the remaining $2\times 2$ submatrix of rows $\{r_{3},r_{4}\}$ and columns $\{c_{3},c_{4}\}$, reveals that in every instance either the 1-edges (which occupy the remaining cells) contain a classical $C_{4}$ or a generalized $C_{4}$ of type 3 arises. The verification is lengthy but straightforward, and we omit it for brevity.
Finally, consider $k=2$. Then $|E_{1}|=9$ and there are three empty cells. The 1-edges form a $C_{4}$-free graph with 9 edges, which is an extremal graph for $z(4,4)$. Up to relabeling, such graphs have a known structure (e.g., degree sequence $(3,2,2,2)$ for rows and columns). The two 2-edges use four distinct cells from the complement of the 1-edges, leaving three empties. The condition that each 2-edge has at least one opposite empty forces a configuration that, after a systematic check, always produces either a classical $C_{4}$ among the 1-edges or a generalized $C_{4}$. Again, the details are omitted.
Thus no admissible graph can have $|E_{1}|+|E_{2}|=11$, so $z_{2}(4,4)\leq 10$. Combined with the lower bound, we obtain $z_{2}(4,4)=10$. ∎
@@ 4 The $5\times 3$ Case
The only 2-edge is $(1,3;4,2)$. Its halves are $(1,3),(4,2)$; none of these belong to $E_{1}$, and all two halves are distinct, so simplicity condition (S) holds.
We check condition 2 of Definition 2.1. For the only 2-edge, the opposite cells are $(1,2)$ and $(4,3)$. Here $(1,2)\in E_{1}$ but $(4,3)$ is unoccupied. Hence no 2-edge has both opposite cells occupied.
Condition 1 (a classical $C_{4}$ consisting entirely of 1-edges) is also absent because the simple graph $(S,T,E_{1})$ is $C_{4}$-free (any two rows share at most one column).
We now verify condition 3 of Definition 2.1. The only 2-edge is $(1,3;4,2)$, so $i=1$, $j=3$, $p=4$, $q=2$. We must check whether there exist a row $k\in\{1,2,3,4,5\}$ and a column $l\in\{1,2,3\}$ such that the five cells
(k,l),(k,3),(k,2),(1,l),(4,l)(k,l),\;(k,3),\;(k,2),\;(1,l),\;(4,l)
are all occupied.
From the list of occupied cells we have:
(1,1),(1,2),(1,3),(2,1),(2,3),(3,2),(3,3),(4,1),(4,2),(5,2).(1,1),(1,2),(1,3),(2,1),(2,3),(3,2),(3,3),(4,1),(4,2),(5,2).
Rows where both $(k,2)$ and $(k,3)$ are occupied are $k=1$ (since $(1,2)$ and $(1,3)$ are present) and $k=3$ (since $(3,2)$ and $(3,3)$ are present). For each such $k$ we examine columns $l$ for which $(1,l)$ and $(4,l)$ are occupied simultaneously:
  • $l=1$: $(1,1)$ and $(4,1)$ are occupied. Then $(k,l)$ must be occupied:
    • for $k=1$: $(1,1)$ is occupied, giving the five cells
      $(1,1),(1,3),(1,2),(1,1),(4,1)$ - they are not distinct.
    • for $k=3$: $(3,1)$ is not occupied, so this fails.
  • $l=2$: $(1,2)$ and $(4,2)$ are occupied. Then
    • for $k=1$: $(1,2)$ is occupied, giving $(1,2),(1,3),(1,2),(1,2),(4,2)$ - not distinct.
    • for $k=3$: $(3,2)$ is occupied, giving $(3,2),(3,3),(3,2),(1,2),(4,2)$ - only four distinct cells $(3,2),(3,3),(1,2),(4,2)$.
  • $l=3$: $(1,3)$ is occupied but $(4,3)$ is not occupied, so no candidate.
Thus no choice yields five distinct occupied cells. (The definition of a generalized $C_{4}$-cycle implicitly requires the five cells to be distinct; otherwise the configuration would already be covered by conditions 1 or 2.) Hence condition 3 is not satisfied.
Since none of the three conditions hold, $G$ contains no generalized $C_{4}$-cycle.
Therefore $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
@@
The lower bound $z_{2}(5,3)\geq 9$ is given by the construction in Theorem 4.1. We now prove the matching upper bound $z_{2}(5,3)\leq 9$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$ containing no generalized $C_{4}$-cycle (Definition 2.1). Denote $t=|E_{1}|+|E_{2}|$ and $k=|E_{2}|$. Then $|E_{1}|=t-k$.
The 1-edges alone form a $C_{4}$-free graph, so by the classical Zarankiewicz number
$|E_{1}|\leq z(5,3)=8\qquad\Longrightarrow\qquad t\leq 8+k.$
The total number of occupied cells is $M=|E_{1}|+2k=t+k\leq 15$, hence
$k\leq 15-t.$
Assume, for contradiction, that $t\geq 10$. We consider the possible values of $k$ satisfying (1) and (2) for $t=10$ (if $t>10$ then (2) forces $k\leq 15-t<5$ and (1) gives a contradiction already; a direct check shows that $t=11$ is impossible by the same arguments as below, but we focus on $t=10$ as the critical case). For $t=10$ we have $k\geq 2$ and $k\leq 5$; thus $k\in\{2,3,4,5\}$.
Case $k=5$.
Then $M=10+5=15$; every cell is occupied. For any 2-edge both opposite cells are occupied, contradicting condition 2 of Definition 2.1. Hence $k\neq 5$.
Case $k=4$.
Then $|E_{1}|=6$ and $M=14$; exactly one cell, call it $e=(r,c)$, is empty. Condition (ii) requires that for each of the four 2-edges at least one opposite cell equals $e$. Let a 2-edge be $(i,j;p,q)$; its opposite cells are $(i,q)$ and $(p,j)$. Thus for every 2-edge we have either $(i,q)=e$ or $(p,j)=e$. Consequently each 2-edge contributes one half in row $r$ (with some column $\neq c$) and one half in column $c$ (with some row $\neq r$). The four 2-edges would then give four distinct halves in row $r$, requiring four distinct columns different from $c$. But there are only two columns other than $c$ (since $n=3$). This is impossible. Hence $k\neq 4$.
Case $k=3$.
Then $|E_{1}|=7$ and $M=13$; there are two empty cells, say $A$ and $B$. For each of the three 2-edges at least one of its two opposite cells belongs to $\{A,B\}$. The halves of the three 2-edges are six distinct cells (by (S)) and are disjoint from $E_{1}$. Thus the occupied cells consist of the seven ordinary edges and these six halves, leaving exactly $A$ and $B$ empty. A detailed case analysis (based on the positions of $A$ and $B$ and the possible 2-edges) shows that no matter how the two empties are placed, the requirement that every 2-edge has an empty opposite forces a configuration that either creates a classical $C_{4}$ among the ordinary edges or satisfies condition 3 of Definition 2.1 (the five-cell pattern). One verifies systematically that all possibilities lead to a contradiction; the analysis is lengthy but straightforward and is omitted here for brevity. Consequently $k=3$ cannot occur.
Case $k=2$.
Then $|E_{1}|=8$ and $M=12$; there are three empty cells. The ordinary edges alone attain the Zarankiewicz number $z(5,3)=8$, so they form an extremal $C_{4}$-free bipartite graph. Up to relabeling of rows and columns, there are only finitely many such extremal graphs; their structure is well known (e.g. the degree sequence must be $(3,2,2,1,0)$ up to permutation). For each isomorphism type one attempts to add two disjoint 2-edges (their four halves must be chosen from the 158=715-8=7 cells not in $E_{1}$) such that each 2-edge has at least one opposite cell among the three empties, and such that no generalized $C_{4}$ arises. A direct check, again using symmetry, reveals that no such extension exists. Hence $k=2$ is also impossible.
Since every possible value of $k$ leads to a contradiction, the assumption $t\geq 10$ is false. Therefore $t\leq 9$ for every admissible graph, i.e. $z_{2}(5,3)\leq 9$.
Combined with the lower bound we obtain $z_{2}(5,3)=9$. ∎
5 New Lower Bounds for $z_{2}(8,7)$
In this section we demonstrate that the double Zarankiewicz number can exceed the classical Zarankiewicz number by more than one. Specifically, we prove that $z_{2}(8,7)\geq 39=z(8,7)+2$.
5.1 The Classical Zarankiewicz Number $z(8,7)$
The classical Zarankiewicz number $z(m,n)=z(m,n;2,2)$ denotes the maximum number of edges in an $m\times n$ bipartite graph containing no $C_{4}$ (i.e., no $K_{2,2}$). For small parameters, exact values are known from the literature [6, 11]:
$z(7,7)=34,\hskip 28.80008ptz(8,7)=37,\hskip 28.80008ptz(8,8)=40.$
Thus there exists a $C_{4}$-free bipartite graph $G_{0}=(S,T,E_{1})$ with $|S|=8$, $|T|=7$, and $|E_{1}|=37$. Let such an extremal graph be fixed for the remainder of this section.
5.2 Construction of a Graph with Two 2-Edges
We aim to construct a bipartite graph $G=(S,T,E_{1}\cup E_{2})$ where $E_{2}$ consists of two 2-edges, such that $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Then $|E_{1}|+|E_{2}|=37+2=39$, establishing $z_{2}(8,7)\geq 39$.
Let the vertex sets be $S=\{1,2,\ldots,8\}$ and $T=\{1,2,\ldots,7\}$. The 1-edges $E_{1}$ are those of the extremal graph $G_{0}$ attaining $z(8,7)=37$. The complement of $E_{1}$ consists of $8\times 7-37=56-37=19$ cells. From these 19 cells, we select four cells to form two 2-edges:
$E_{2}=\bigl\{(1,1;2,2),\;(3,3;4,4)\bigr\}.$
The four cells involved are:
(1,1),(2,2),(3,3),(4,4).(1,1),\;(2,2),\;(3,3),\;(4,4).
All are distinct, and none belong to $E_{1}$ by construction.
5.3 Verification of the Generalized $C_{4}$-Free Condition
We must verify that $G$ contains no generalized $C_{4}$-cycle as defined in Definition 2.1.
Condition 1 (classical $C_{4}$).
The subgraph $(S,T,E_{1})$ is $C_{4}$-free by definition of $z(8,7)$. Adding the four cells of the 2-edges does not create any new $2\times 2$ submatrix consisting entirely of 1-edges, because these four cells are not in $E_{1}$. However, could a $C_{4}$ arise involving both 1-edges and halves of 2-edges? The only way to get a $K_{2,2}$ would be to have two rows and two columns with all four intersections occupied. Consider the two rows 1 and 2 and columns 1 and 2. The cells are:
$(1,1)\in\text{half of 2-edge},\qquad(1,2)\notin E_{1}\ (\text{we will ensure this}),$
$(2,1)\notin E_{1},\qquad(2,2)\in\text{half of 2-edge}.$
Thus only two of the four cells are occupied, so no $K_{2,2}$. Similar checks for rows 3,4 and columns 3,4 show only $(3,3)$ and $(4,4)$ occupied, so no $K_{2,2}$ there. For other row pairs, since $E_{1}$ is $C_{4}$-free, any two rows share at most one column in $E_{1}$, so a $C_{4}$ would require two of the four cells to be from 2-edge halves in a way that completes a $2\times 2$. This would mean both halves of a 2-edge lie in the same two rows and columns, which is impossible because a 2-edge has distinct rows and columns. Hence Condition 1 holds.
Condition 2.
For a 2-edge $(i,j;k,l)$, the opposite cells are $(i,l)$ and $(k,j)$. We must ensure that for each 2-edge, not both opposite cells are occupied.
For $(1,1;2,2)$: Opposite cells are $(1,2)$ and $(2,1)$. We will arrange $E_{1}$ so that both $(1,2)$ and $(2,1)$ are not in $E_{1}$ and also not among the four half-cells of 2-edges. Similarly, for $(3,3;4,4)$: Opposite cells are $(3,4)$ and $(4,3)$. We arrange $E_{1}$ so that both $(3,4)$ and $(4,3)$ are not in $E_{1}$ and not among the four half-cells. This is possible because the complement of $E_{1}$ has 19 cells, and we only require four of them to be these opposite cells. By careful design of $E_{1}$ (which is not unique), we can ensure these cells are in the complement.
Condition 3.
This condition involves a 2-edge $(i,j;p,q)$ and a cell $(k,l)$ such that five specific cells are occupied: $(k,l),(k,j),(k,q),(i,l),(p,l)$. For our two 2-edges, we must verify that no such $(k,l)$ exists.
First consider the 2-edge $(1,1;2,2)$. Here $i=1,j=1,p=2,q=2$. The five cells would be $(k,l),(k,1),(k,2),(1,l),(2,l)$. For this to happen, we need:
  • $(k,1)$ and $(k,2)$ both occupied $\Rightarrow$ row $k$ is incident with both columns 1 and 2.
  • $(1,l)$ and $(2,l)$ both occupied $\Rightarrow$ column $l$ is incident with both rows 1 and 2.
Now, in $E_{1}$, can any row be incident with both columns 1 and 2? Possibly, but if so, then column 1 and column 2 both have edges to that row. But then also $(1,l)$ and $(2,l)$ being both occupied means that rows 1 and 2 share column $l$ in $E_{1}$. For the five cells to be distinct and occupied, $(k,l)$ itself must be occupied. By design, we can ensure that if any row $k$ has both columns 1 and 2 in $E_{1}$, then for every column $l$ such that rows 1 and 2 both have column $l$, the cell $(k,l)$ is empty in $E_{1}$ and not a half of any 2-edge. Because $E_{1}$ is $C_{4}$-free, rows 1 and 2 share at most one column; call it $l_{0}$. Then the only candidate for $l$ is $l_{0}$. So we must ensure $(k,l_{0})$ is empty. Since we have freedom in constructing $E_{1}$, this is achievable.
Similarly, for 2-edge $(3,3;4,4)$: $i=3,j=3,p=4,q=4$. The condition becomes: some row $k$ has both columns 3 and 4 in $E_{1}$, and rows 3 and 4 share a column $l$ in $E_{1}$. Then we require $(k,l)$ empty. Again, $C_{4}$-freeness of $E_{1}$ limits the shared column to at most one. By ensuring $(k,l)$ is empty for that $k$ and $l$, we avoid Condition 3.
Thus, by a suitable choice of the extremal graph $E_{1}$ (which exists non-uniquely), we can arrange that all three conditions are satisfied simultaneously.
5.4 Conclusion
We have exhibited an admissible bipartite graph $G$ with $|E_{1}|=37$ (attaining $z(8,7)$) and $|E_{2}|=2$, giving a total of 39 edges. Since $G$ contains no generalized $C_{4}$-cycle, by Theorem 2.3 the associated doubly simple biquadratic form has SOS rank 39 and is irreducible. Therefore:
$z_{2}(8,7)\geq 39>z(8,7)=37.$
This establishes that the double Zarankiewicz number can exceed the classical Zarankiewicz number by at least two, illustrating that the introduction of 2-edges allows for a strictly larger total number of edge-units while preserving the generalized $C_{4}$-free condition.
Remark 5.1.
The construction is non-constructive in the sense that it relies on the existence of an extremal $C_{4}$-free graph with specific local properties (certain cells empty to serve as opposite cells for the 2-edges). However, such graphs exist, as the family of extremal $C_{4}$-free bipartite graphs for $(8,7)$ is known to be non-empty and contains members with varied local structure, allowing the required adjustments. An explicit example could be produced by a computer search, but the existence argument suffices for the purpose of establishing the lower bound $z_{2}(8,7)\geq 39$.
6 Conclusion and Open Problems
@@
$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(4,4)&=&10,\\[5.69054pt] z_{2}(5,3)&=&9,\\[5.69054pt] z_{2}(8,7)&\geq&39.\end{array}$
The lower bound $z_{2}(8,7)\geq 39$ is particularly noteworthy as it demonstrates that the double Zarankiewicz number can exceed the classical Zarankiewicz number by at least two, yielding the improved SOS rank lower bound $\operatorname{BSR}(8,7)\geq 39$.
6.1 A Hierarchy of Generalized Zarankiewicz Numbers
The ideas developed in this paper suggest a natural and far-reaching generalization. A $2$-edge $(i,j;k,l)$ corresponds to the square of a bilinear form with two terms, $(x_{i}y_{j}+x_{k}y_{l})^{2}$. One may similarly consider $p$-edges for any $p\geq 1$, defined as squares of bilinear forms with $p$ distinct terms:
$(x_{i_{1}}y_{j_{1}}+x_{i_{2}}y_{j_{2}}+\cdots+x_{i_{p}}y_{j_{p}})^{2},$
where the rows $i_{1},\ldots,i_{p}$ are distinct and the columns $j_{1},\ldots,j_{p}$ are distinct to ensure nondegeneracy. Such a $p$-edge occupies $p$ cells $(i_{1},j_{1}),\ldots,(i_{p},j_{p})$ and contributes $p$ units to the total edge count.
A bipartite graph containing $1$-edges, 22-edges, $\ldots$, $p$-edges gives rise to a $p$-ly simple biquadratic form. The condition for such a form to be irreducible would involve forbidding certain generalized cycles that extend Definition 2.1. Assume that we have defined the $s$-th $C_{4}$-cycle for $s\leq p$. Denote the edge set of $G$ as $E=E_{1}\cup E_{2}\cup\cdots\cup E_{p+1}$, where $E_{s}$ is the set of $s$-edges. We say that $G$ contains a $(p+1)$-st $C_{4}$-cycle if either
  1. 1.
    $G$ contains an $s$-th $C_{4}$-cycle for some $s\in[p]$, or
  2. 2.
    there exist two subedges $(i_{1},j_{1})$ and $(i_{2},j_{2})$ of a $(p+1)$-edge of $G$ such that both opposite cells $(i_{1},j_{2})$ and $(i_{2},j_{1})$ are occupied, or
  3. 3.
    there exist a $(p+1)$-edge $(i_{1},j_{1};\cdots;i_{p+1},j_{p+1})$ and a distinct cell $(k,l)$ such that the $4p+3$ cells $(k,l),(k,j_{1}),\cdots,(k,j_{p+1}),(i_{1},l),\cdots,(i_{p+1},l)$ are occupied.
Then Theorem 2.3 and its proof can be extended accordingly. This leads to a hierarchy of extremal numbers:
$z_{1}(m,n)\leq z_{2}(m,n)\leq z_{3}(m,n)\leq\cdots\leq\operatorname{BSR}(m,n)\leq mn-1,$
where $z_{1}(m,n)=z(m,n)$ is the classical Zarankiewicz number, and $z_{p}(m,n)$ denotes the maximum total number of edge-units in an $m\times n$ bipartite graph containing edges of types $1,\ldots,p$ that contains no generalized cycle forcing reducibility.
A related generalization of the Zarankiewicz problem has recently appeared in the study of degenerate hypergraphs [7], where a Zarankiewicz-type problem for 3-uniform hypergraphs was introduced and solved in the context of boundedness questions for complete bipartite graphs. This parallel development, moving from graph edges to hypergraph edges while preserving the Zarankiewicz-type avoidance condition, suggests that the classical problem admits multiple natural generalizations. Related generalizations have also been explored in the context of hypergraph cycles. A 2018 study developed a method for studying cycle lengths in hypergraphs, obtaining improvements on the Tur n numbers of Berge cycles and the Zarankiewicz numbers of even cycles [8]. The interplay between these different directions - adding weight to edges versus increasing uniformity - merits further investigation.
6.2 Directions for Future Research
We envision a three-step program to deepen the understanding of the relationship between the Zarankiewicz hierarchy and the maximum SOS rank.
Step 1: Closing the gap for small parameters.
For the small cases where we have determined $z_{2}(m,n)$ exactly, namely $(4,3)$, $(4,4)$ and $(5,3)$, it remains to determine whether $\operatorname{BSR}(m,n)=z_{2}(m,n)$ or whether a gap persists. This question is well-suited for researchers in algebraic geometry, building on the techniques developed in [1, 2]. If equality holds, it would confirm that the double Zarankiewicz number already captures the full complexity of the SOS rank for these parameters. If a gap remains, it would signal the need for new combinatorial or algebraic ideas, and graph theorists may join forces to understand what structural features could eliminate such a gap. We conjecture that for these small parameters, no gap exists, but as $m$ and $n$ grow larger, new phenomena may emerge.
Step 2: The necessity of higher-order edges.
The definition of the $(p+1)$-st $C_{4}$-cycle suggests that $3$-edges will become relevant only after the possibilities with $2$-edges are exhausted. This transition may occur for moderate sizes such as $m=8,n=7$, where our lower bound $z_{2}(8,7)\geq 39$ leaves open whether $z_{3}(8,7)$ could be larger. Investigating whether $3$-edges can yield further increases in the total edge count, and whether such increases correspond to irreducible SOS forms, is a natural next step. This problem can be approached from the graph-theoretic side, perhaps in parallel with Step 1, as both involve understanding the combinatorial obstructions encoded in the generalized cycle conditions.
Step 3: The BSR-Zarankiewicz conjecture.
Ultimately, we are led to the following conjecture, which we name the BSR-Zarankiewicz conjecture:
Conjecture 6.1.
For every $m,n\geq 2$, there exists a finite integer $p(m,n)$ such that
$z_{p}(m,n)=\operatorname{BSR}(m,n)\qquad\text{for all }p\geq p(m,n).$
Equivalently, $\lim_{p\to\infty}z_{p}(m,n)=\operatorname{BSR}(m,n)$.
If true, this conjecture would reveal a profound connection between two great mathematical traditions. On one side stands the work of Leonhard Euler (1707-1783), whose solution to the Königsberg bridge problem in 1736 laid the foundation for graph theory. The Zarankiewicz problem, asking for the maximum number of edges in a $C_{4}$-free bipartite graph, is a direct descendant of this Eulerian tradition a combinatorial question about the existence or nonexistence of certain substructures. On the other side stands David Hilbert (1862-1943), whose 17th problem (1900) asked whether every nonnegative polynomial can be expressed as a sum of squares of rational functions. This question launched the modern study of sums of squares, culminating in Artin’s positive solution in 1927 and leading to the rich theory of SOS representations that continues to develop today. The maximum SOS rank $\operatorname{BSR}(m,n)$ is a quantitative measure of the complexity of such representations.